OpenStudy (anonymous):

8/3x+6 Added by 6x/x^2 +x -2

OpenStudy (aum):

\( \Large \frac{8}{3x+6}+\frac{6x}{x^2+x-2} = \frac{8}{3(x+2)}+\frac{6x}{(x+2)(x-1)}\\ \Large \) Find the least common denominator and add the fractions. Simplify.